# Activated solids: Spontaneous deformations, non-affine fluctuations, softening, and failure

### v1

@@ IV.4 Topological defects
We use Voronoi tessellations to identify each particle’s topological neighbors [81]. In a perfect lattice, each particle has 6 neighbors, with 5- or 7-fold deviations indicating topological defects. Minimal fluctuations in solid accommodate bound 5–7–5–7 dislocation pairs. The quilibrium BKTHNY theory links solid-hexatic melting to the unbinding of these pairs, creating dislocations, with hexatic melting driven by dislocation-unbinding to free 5- or 7-fold disclinations.
@@
V Local control of non-affinity and defect
Controlling local mechanical properties in solids has long been a challenge in physics and material science [82]. This can be achieved by manipulating non-affinity and defects, especially in colloidal crystals [83]. Previous studies in equilibrium solids used complex local laser fields [84] to drive non-affine parameters [39, 85]. Building on our study of non-affinity in activated melting, we propose a simpler method for local mechanical control: applying laser power to self-thermophoretic Janus colloids [64], inducing non-affinity and/or defects in an otherwise equilibrium solid. Localized activity control can also be achieved in other active colloidal or robotic systems [20, 86].
Figure 8: (a) Schematic of the simulation setup, where particles are selectively activated within a circular region of radius $r_{\rm beam}=30\sigma$ at the center of the simulation box, simulating laser irradiation. (b), (c) Snapshots of the system under $\Lambda=11$ at times $t=2\times 10^{6}\tau_{p}$ and $6.02\times 10^{6}\tau_{p}$, with particles colored by their local non-affinity $\chi_{i}$ and defects (particles with non-six neighbors) indicated. (d)–(f) Time evolution of the global non-affine deviation $\Delta X$ inside (blue) and outside (orange) the active region. (g)–(i) Evolution of defect fraction inside (blue) and outside (orange) the active region. Parameter values are $\Lambda=9$ (d,g), $\Lambda=10$ (e,h), and $\Lambda=11$ (f,i).
To study the above-mentioned effect near the solid–hexatic transition, we examine a collection of selectively activated particles coupled to a thermal heat bath with $\tilde{D}_{t}=1$, activated over a central circular region (Fig.8). Three activity levels are considered: below ($\Lambda=9$), at ($\Lambda=10$), and above ($\Lambda=11$) the transition. The system consists of $N=16384$ particles, initially arranged in a triangular lattice at density $\rho=1.1$ within a rectangular simulation box of size $131.14\sigma\times 113.57\sigma$. After sufficient equilibration, the central circular region of radius $r_{\rm beam}=30\sigma$ is selectively activated to mimic laser irradiation (Fig.8(a)). Particles are assigned activity $\Lambda_{i}(t)=\sigma_{i}(t)\Lambda$, with $\sigma_{i}=1$ inside the active zone and $\sigma_{i}=0$ outside. The value of $\sigma_{i}(t)$ is updated at each timestep based on particle positions.
@@ V Local control of non-affinity and defect
A schematic of this setup is shown in Fig. 8(a), and representative snapshots during evolution for $\Lambda=11$ are shown in Fig. 8(b), (c), where particles are colored by their local non-affinity $\chi_{i}$. Pronounced softening and defect formation are observed inside the active region, while $\chi_{i}$ outside remains close to equilibrium. At early times, defects remain confined within the active zone; however, at later times, isolated defects begin to appear in the passive region as a result of migration and delayed annihilation of defects originating from the active area. To quantify global structural changes, we define $\Delta X=X(\Lambda)-X(0)$, the deviation of the global non-affine parameter from its equilibrium value, as defined in Sec. III.2. The evolution of this quantity in the active zone $\Delta X^{\rm(in)}$ and in the passive zone $\Delta X^{\rm(out)}$ is shown in Fig. 8(d)–(f). Below the transition ($\Lambda=9$), $\Delta X^{\rm(in)}$ fluctuates around a non-zero mean with a larger amplitude than in equilibrium, while $\Delta X^{\rm(out)}$ remains negligible. At transition ($\Lambda=10$), $\Delta X^{\rm(in)}$ diverges, while $\Delta X^{\rm(out)}$ stays flat, indicating localized melting. Above transition ($\Lambda=11$), this divergence occurs earlier followed by an increase in $\Delta X^{\rm(out)}$ as well, suggesting outward spread of non-affineness. The defect fraction evolution in Fig. 8(g)-(i) shows that for $\Lambda=9$, defects form and annihilate intermittently, while at and beyond the transition, defect fraction increases significantly inside and, to a lesser extent, outside the active region.
@@ VI Discussion and Conclusion
In summary, this paper explored spontaneous deformations in active solids, characterizing crystal deformations through non-affine fluctuations and deriving a scaling law in terms of activity $\Lambda$ and density. Our scaling analysis and numerical data collapse predict that non-affinity scales as $\Lambda^{2}$ with activity and $(\rho-\rho_{m})^{-1}$ with density. At large $\Lambda$, the non-affine parameter diverges, signaling defect proliferation and the activated melting to a hexatic phase with vanishing shear modulus and order. This softening was further analyzed through topological defect proliferation. At higher activity, the hexatic melts to a fluid as captured by the change in structure factor and hexatic order parameter. Using the properties found from our studies, we finally proposed, locally tuned activity as a means to control non-affinity and mechanical properties. We numerically demonstrated the use of an isotropic activity patch in the solid. Such a patch locally softens the solid by increasing non-affinity, which reduces the shear modulus. Activating the solid in an anisotropic domain can modify elastic properties differently in each direction. Our predictions can be directly tested in active colloids. Additionally, locally controlling non-affinity via activity is simpler to implement in Janus colloid experiments than previous laser-controlled methods [39].
@@
The work began with a discussion between DC and SS, followed by initial non-affinity scaling simulations carried out by DD. PN set up and conducted all the final numerical simulations. DC supervised the project, and all authors contributed to writing the manuscript and discussed the results.
@@ Acknowledgements.
DC acknowledges research grants from the Department of Atomic Energy (OM no. 1603/2/2020/IoP/R&D-II/15028), an Associateship at ICTS-TIFR, Bangalore, and expresses gratitude to MPIPKS, Dresden for hospitality during a two-month visit in 2024, where part of the research was carried out. The simulations were performed using resources at IOP Bhubaneswar. DC, DD, and PN sincerely thank Surajit Sengupta [87], who tragically passed away, for his advocacy of using non-affine fluctuations to characterize active solids, and dedicate this paper to his memory.
@@ Appendix A Dimensionless equations
To perform numerical simulations, we use the dimensionless equations by applying the following substitutions: ${\bf r}_{i}\to\tilde{\bf r}_{i}={\bf r}_{i}/\sigma$, $t\to\tilde{t}=t/\tau_{u}=t\mu\epsilon/\sigma^{2}$, the dimensionless activity $\Lambda=v_{0}\tau_{u}/\sigma=v_{0}\sigma/\mu\epsilon$, the dimensionless translational diffusivity $D_{t}\to\tilde{D}_{t}=D_{t}\tau_{u}/\sigma^{2}=D_{t}/\mu\epsilon$, $\nabla_{i}\to\tilde{\nabla}_{i}=\sigma\nabla_{i}$, $U(r_{ij})\to\tilde{U}=U/\epsilon$, $\tilde{\mu}=\mu\epsilon\tau_{u}/\sigma^{2}=1$ and the dimensionless rotational diffusivity $\tilde{D}_{r}=D_{r}\sigma^{2}/\mu\epsilon$. By using these substitutions, one obtains the dimensionless form of the equations
$d\tilde{\bf r}_{i}(\tilde{t})$ $=\Lambda\,\hat{\bf n}_{i}(\tilde{t})d\tilde{t}-\tilde{\nabla}_{i}\sum_{j\epsilon R_{i}}\tilde{U}(r_{ij})d\tilde{t}+\sqrt{2\tilde{D}_{t}d\tilde{t}}\,d{\bf B}_{i}(\tilde{t})$
$d{\theta_{i}}(\tilde{t})$ $=(2\,\tilde{D}_{r}d\tilde{t})^{1/2}\,dB^{r}_{i}(\tilde{t}).$ (13)
We perform numerical simulations with $\tilde{D}_{r}=3$, $\tilde{D}_{t}=1$ for thermal, and $\tilde{D}_{t}=0$ for athermal cases.
@@ Appendix B Scaling ansatz for $X$ at different system sizes
Figure 9: As in Fig. 2, we plot $\langle\Delta X\rangle$ for two system sizes ($N=4096,16384$) at density $\rho\sigma^{2}=1.1$, demonstrating that the scaling $\langle\Delta X\rangle\sim\Lambda^{2}$ holds for both the thermal ($\tilde{D}_{t}=1$) and athermal ($\tilde{D}_{t}=0$) cases at the larger system size.
@@
In Fig. 2 of the main text, we showed the scaling behavior $\langle\Delta X\rangle\sim\Lambda^{2}$ for a system of size $N=4096$. To verify the robustness of this scaling, we present additional data for a larger system size ($N=16384$) at density $\rho\sigma^{2}=1.1$ in Fig. 9. We confirm that the quadratic scaling $\langle\Delta X\rangle\sim\Lambda^{2}$ remains valid for both thermal ($\tilde{D}_{t}=1$) and athermal ($\tilde{D}_{t}=0$) cases, consistently across system sizes.
Appendix C Obtaining shear modulus
@@
Figure 10: (a) Shear stress, $\sigma_{xy}^{tot}$, vs. shear strain, $\mathcal{E}_{xy}=\gamma$, at $\Lambda=7$ for 24 ensembles without significant non-linearity or hysteresis. Different colors represent individual ensembles, with the black line showing the linear fit, where the slope gives the shear modulus, $\langle{\cal G}\rangle=105.7\pm 0.679$ in units of $\epsilon/\sigma^{2}$. (b) Linear response trajectories and (c) non-linear response trajectories exhibiting significant hysteresis and nonlinearity. Non-linear trajectories are excluded from shear modulus calculation. Raw data are shown as semi-transparent circular markers, with opaque square markers representing averages over 100 consecutive values of $\gamma$ and $\sigma^{\mathrm{tot}}_{xy}$. Markers are color-coded by shear direction: forward shear $0\rightarrow\gamma_{m}$ in red, backward shear $\gamma_{m}\rightarrow-\gamma_{m}$ in green, and $-\gamma_{m}\rightarrow 0$ in blue. (d) Fraction of linear ensembles, $n_{\mathrm{lin}}/n_{\mathrm{tot}}$, as a function of activity $\Lambda$.
@@
[…]
at a strain rate ˙$\dot{{\cal E}}=10^{-2}D_{r}$ and ${\cal E}_{xy}^{\max}=0.035$. The shear modulus is a linear response coefficient; thus the chosen ${\cal E}_{xy}^{\max}$ is to ensure the system response stays within the linear regime for accurate shear modulus extraction.
For $\Lambda\leq 6$, all four initial configurations exhibit linear stress-strain behavior. For $\Lambda>6$, non-linear deviations appear. We generate $n_{\mathrm{tot}}\approx 100$ initial configurations, discarding those with nonlinearity for shear modulus analysis, which becomes more common at higher $\Lambda$. The shear modulus is then computed by averaging the linear response data from the remaining $n_{\mathrm{lin}}$ hysteresis-free ensembles, as shown in Fig. 10(a). Examples of linear and non-linear ensembles are shown in Figs. 10(b) and (c).

### v2

@@ IV.4 Topological defects
We use Voronoi tessellations to identify each particle’s topological neighbors [43]. In a perfect lattice, each particle has 6 neighbors, with 5- or 7-fold deviations indicating topological defects. Minimal fluctuations in solid accommodate bound 5–7–5–7 dislocation pairs. According to equilibrium BKTHNY theory, solid-hexatic melting occurs via the unbinding of these dislocation pairs, creating free dislocations, while the transition from hexatic to liquid is driven by the unbinding of dislocations into free 5- or 7-fold disclinations. Since a detailed characterization of BKTHNY-like melting in active solids has been presented previously [20, 39], we do not revisit measurements of order-parameter correlations or scaling exponents here.
In Fig. 9(a),(b) we identify such defects in typical configurations across activated solid melting. We identify clusters of such defects, including five, seven, and eight fold defects. We use fraction of particles with $\nu$ neighbors, $n_{\nu}=\langle N_{\nu}\rangle/N$ to analyze the defect formation. Here $N_{\nu}$ indicates the number of particle with $\nu=5,6,7$ neighbors and $\langle\dots\rangle$ denotes the steady state average. Although larger departures from $\nu=6$ are present, their fraction is significantly low.
V Local softening: control of non-affinity and defect
Controlling local mechanical properties in solids is a longstanding challenge in physics and materials science [44], particularly for colloidal crystals where structure and elasticity can be tuned at the particle scale [45, 46]. Earlier attempts to control non-affine fluctuations in passive solids proposed introducing a force field conjugate to non-affine modes as a theoretical route to control local non-affinity [41, 47]. This field corresponds to an effective many-body force that depends explicitly on the instantaneous particle configuration within a local neighborhood, with fixed reference lattice positions entering as parameters. The implementation scheme based on dynamically modulated laser traps outlined in [41], therefore, requires real-time particle tracking and the synthesis of highly structured, configuration-dependent forces, making experimental realization extremely challenging.
In the present work, building on our analysis of activity-induced non-affine fluctuations, we propose a qualitatively distinct and experimentally accessible strategy for controlling non-affinity through locally tunable activity. Specifically, optically activated self-thermophoretic Janus colloids [48], combined with a simple defocused laser beam, can be used to locally activate a region of an otherwise passive colloidal crystal, thereby robustly inducing non-affine deformations. Similar localized activation protocols are feasible in a broad class of active colloidal and robotic systems [18, 49]. This approach enables direct spatial modulation of the activity strength $\Lambda$, which in turn locally controls the degree of non-affinity according to $\langle X\rangle\sim\Lambda^{2}$. Crucially, this approach circumvents the need for real-time configuration tracking or the synthesis of complex many-body forces, and instead relies on the system’s intrinsic active dynamics to achieve local, tunable control of non-affine fluctuations.
To investigate these effects near the solid–hexatic transition, we simulate a large system of $N=16384$ particles arranged initially in a triangular lattice at density $\rho\sigma^{2}=1.1$, coupled to a thermal heat bath with diffusivity $\tilde{D}_{t}=1$, and a fixed $\tilde{D}_{r}=3.0$. The system evolves in a rectangular simulation box of size $131.14\sigma\times 113.57\sigma$, under periodic boundary conditions. After sufficient equilibration, activity is applied selectively within a central circular region (radius $r_{\rm beam}=30\sigma$), as illustrated in Fig. 10(a). We explore three activity levels: below ($\Lambda=9$), at ($\Lambda=10$), and above ($\Lambda=11$) the transition threshold. The activity of each particle $\Lambda_{i}(t)$ is implemented via a binary field $\sigma_{i}(t)$, such that:
$\Lambda_{i}(t)=\sigma_{i}(t)\Lambda,$
@@ V Local softening: control of non-affinity and defect
A schematic of this setup is shown in Fig. 10(a). Snapshots for $\Lambda=11$ in Fig. 10(b,c) show evolution over time, where particles are colored by their local non-affinity $\chi_{i}$, display pronounced softening and defect formation confined within the activated zone, while $\chi_{i}$ remains near equilibrium outside. At early times, defects remain localized in the active zone, but later, isolated defects appear in the passive region due to migration and delayed annihilation of those originating from the active zone. To quantify structural changes, we define $X^{(a)}=X(\Lambda)-X(0)$, the shift in global non-affinity from equilibrium, as defined in Sec. III.2. The evolution of this quantity in the active zone $X^{(a)}_{\rm in}$ and in the passive zone $X^{(a)}_{\rm out}$ is shown in Fig. 10(d)–(f). These figures show that below the transition, $X^{(a)}_{\rm in}$ fluctuates around a steady mean with enhanced amplitude, while $X^{(a)}_{\rm out}$ stays negligible. At the transition, $X^{(a)}_{\rm in}$ diverges, with no corresponding rise in $X^{(a)}_{\rm out}$, indicating localized melting. Above the transition, divergence in $X^{(a)}_{\rm in}$ occurs earlier and is followed by an increase in $X^{(a)}_{\rm out}$, suggesting outward propagation of non-affineness.
The evolution of defect fraction in Fig. 10 (g-i) complements these observations. For $\Lambda=9$, defects intermittently form and annihilate; for $\Lambda\geq 10$, the defect fraction increases significantly within the activated region and more gradually outside it. The simultaneous rise of non-affinity and defect fraction suggests a coupled onset of mechanical disorder at these high activity.
Figure 10: (a) Schematic of the simulation setup, where particles are selectively activated within a circular region of radius $r_{\rm beam}=30\sigma$ at the center of the simulation box, simulating laser irradiation. (b), (c) Snapshots of the system under $\Lambda=11$ at times $t=2\times 10^{6}\tau_{p}$ and $6.02\times 10^{6}\tau_{p}$, with particles colored by their local non-affinity $\chi_{i}$ and defects (particles with non-six neighbors) indicated. (d)–(f) Time evolution of the excess non-affinity $X^{(a)}$ inside (blue) and outside (orange) the active region. (g)–(i) Evolution of defect fraction inside (blue) and outside (orange) the active region. Parameter values are $\Lambda=9$ (d,g), $\Lambda=10$ (e,h), and $\Lambda=11$ (f,i).
@@ VI Discussion and Conclusion
In this work, we explored spontaneous deformations in active solids by quantifying local rearrangements using non-affine fluctuations relative to global strain. Our main achievements are three fold.
First, using scaling arguments and numerical simulations, we obtained precise predictions for the steady-state mean non-affinity as a function of activity parameters, such as active speed and persistence. The system-averaged non-affinity grows quadratically with active speed, scaling as $\Lambda^{2}$. When examined as a function of active persistence time, it shows an initial linear increase followed by saturation – behavior that cannot be captured by simple active-temperature estimates. Additionally, the non-affinity scales inversely with the distance from the melting density, $(\rho-\rho_{m})^{-1}$.
Second, at higher activity, the distribution of local non-affinity reveals the emergence of pronounced non-affine zones coexisting with regions of predominantly small non-affine fluctuations. As activity increases, the distributions broaden, become more skewed, and develop heavy tails, eventually exhibiting a secondary maximum indicative of coexisting low and high non-affinity regions. This coexistence precedes the divergence of non-affinity, which acts as a precursor to defect proliferation and eventual melting. Steady-state spatial correlations of local non-affinity exhibit a correlation length that grows with active speed and increases with persistence before saturating. The dynamical relaxation of system-averaged non-affinity is governed by the active persistence time.
Activity-induced softening is accompanied by a reduction in the shear modulus and the loss of both solid and hexatic order. The solid shows a two-step melting in a manner similar to the BKTHNY scenario, mediated by proliferation of topological defects.
Third, we demonstrated that spatially patterned activation provides a direct route to locally induce non-affinity and mechanical softening. Unlike previous schemes requiring complex, configuration-dependent many-body forces, our approach is experimentally accessible: by selectively activating regions of an otherwise passive colloidal crystal – such as with optically controlled self-thermophoretic Janus particles – non-affine deformations can be reliably induced. This method directly tunes activity to control non-affinity, avoids real-time particle tracking, and leverages the system’s intrinsic active dynamics, offering a practical and experimentally feasible pathway for local, tunable mechanical control.
More generally, anisotropic activation protocols may enable directional control of solid mechanics, opening avenues for spatially programmable mechanical responses. Beyond enriching our understanding of active solid mechanics, these insights contribute broadly to the design of adaptive metamaterials and may inform how biological systems dynamically regulate their rigidity through internally generated stresses.
@@
The work began with a discussion between DC and SS, followed by initial non-affinity scaling simulations carried out by DD. PN set up and conducted all the final numerical simulations and analyzed the results. DC supervised the project and wrote the final manuscript; all authors contributed to initial writing and discussed the results.
@@
Conflicts of interest
There are no conflicts of interest to declare.
@@ Acknowledgements.
DC acknowledges support from the Department of Atomic Energy (OM no. 1603/2/2020/IoP/R&D-II/15028), an Associateship at ICTS-TIFR, Bangalore, and thanks MPIPKS, Dresden, for hospitality during a two-month visit in 2024, where part of this work was done. DC also thanks Tamoghna Das for useful comments. Numerical simulations were performed using the SAMKHYA high-performance computing cluster and other computational resources at the Institute of Physics, Bhubaneswar. This work is dedicated to the memory of our late co-author, Prof. Surajit Sengupta [50], in recognition of his lasting contributions to the field and to this research.
Figure 11: As in Fig. 2, we plot $\langle X^{(a)}\rangle$ for two system sizes ($N=4096,16384$) at density $\rho\sigma^{2}=1.1$, demonstrating that the scaling $\langle X^{(a)}\rangle\sim\Lambda^{2}$ holds for both the thermal ($\tilde{D}_{t}=1$) and athermal ($\tilde{D}_{t}=0$) cases at the larger system size.
@@ Appendix A Dimensionless equations
To perform numerical simulations, we use the dimensionless equations by applying the following substitutions: ${\mathbf{r}}_{i}\to\tilde{\mathbf{r}}_{i}={\mathbf{r}}_{i}/\sigma$, $t\to\tilde{t}=t/\tau_{u}=t\mu\epsilon/\sigma^{2}$, the dimensionless activity $\Lambda=v_{0}\tau_{u}/\sigma=v_{0}\sigma/\mu\epsilon$, the dimensionless translational diffusivity $D_{t}\to\tilde{D}_{t}=D_{t}\tau_{u}/\sigma^{2}=D_{t}/\mu\epsilon$, $\nabla_{i}\to\tilde{\nabla}_{i}=\sigma\nabla_{i}$, $U(r_{ij})\to\tilde{U}=U/\epsilon$, $\tilde{\mu}=\mu\epsilon\tau_{u}/\sigma^{2}=1$ and the dimensionless rotational diffusivity $\tilde{D}_{r}=D_{r}\sigma^{2}/\mu\epsilon$. By using these substitutions, one obtains the dimensionless form of the equations {align} d ~r_i(~t) = Λ  ^n_i(~t) d ~t - ~∇_i∑_j ϵR_i ~U(r_ij) d ~t + 2 ~D_t  dB_i(~t)
dθ_i( ~t) = (2   ~D_r d ~t)^1/2 dB^r_i( ~t).
We perform numerical simulations with $\tilde{D}_{t}=1$ for thermal, and $\tilde{D}_{t}=0$ for athermal cases.
@@ Appendix B Scaling ansatz for $X$ at different system sizes
In Fig. 2 of the main text, we showed the scaling behavior $\langle X^{(a)}\rangle\sim\Lambda^{2}$ for a system of size $N=4096$. To verify the robustness of this scaling, we present additional data for a larger system size ($N=16384$) at density $\rho\sigma^{2}=1.1$ in Fig. 11. We confirm that the quadratic scaling $\langle X^{(a)}\rangle\sim\Lambda^{2}$ remains valid for both thermal ($\tilde{D}_{t}=1$) and athermal ($\tilde{D}_{t}=0$) cases, consistently across system sizes.
Appendix C Probability distributions of $X$
In Figure 12, we explore the probability distributions of the global non-affine parameter $X$ for an ABP solid at $\tilde{D}_{r}=3.0$, $\tilde{D_{t}}=1$ and $\rho\sigma^{2}=1.1$. Figure 12(a) compares $P(X)$ for $N=4096$ particles across different $\Lambda$ values, showing that as $\Lambda$ increases, the mode of the distribution shifts to larger $X$ (consistent with the $\langle X\rangle\sim\Lambda^{2}$ scaling) and broadens as $\langle X^{2}\rangle-\langle X\rangle^{2}\sim\Lambda^{4}$ (plot not shown). This behavior captures softening even in the absence of defect formation (as noted in the main text, the onset of defect formation appears at $\Lambda\approx 10$). As shown in Fig. 12(b) for $\Lambda=4$ and $N$ ranging from 1024 to 16384, the probability distribution $P(X)$ is nearly Gaussian with variance $\sigma^{2}\sim 1/N$, consistent with the central limit theorem for the mean of uncorrelated entities. The vanishing variance of $X$ with $N$ characterizes the intensiveness of the global non-affine parameter, characterizing the particle rearrangements expelling local stress. Similar scaling behavior for fluctuations in $X$ was observed before in equilibrium solids [41].
Figure 13: Stacked probability distributions $P(\chi)$ of local non-affinity on log-log axes for varying activity $\Lambda$ (rows) and rotational diffusion $\tilde{D}_{r}$ (columns). Increasing $\Lambda$ shifts and broadens the distributions, while increasing $\tilde{D}_{r}$ suppresses large-$\chi$ tails; bimodality emerges at large $\Lambda$ for $\tilde{D}_{r}<0.25$.
Appendix D Probability distributions of local non-affinity
Figure 13 compiles stacked probability distributions $P(\chi)$ of the local non-affinity $\chi$ on log-log axes over a broad range of active speed $\Lambda$ and rotational diffusivity of the heading direction $\tilde{D}_{r}$. Rows correspond to increasing activity, $\Lambda=1,4,5,6,7$, while columns span rotational diffusivities $\tilde{D}_{r}=0.010,0.025,0.100,0.250,1,3,10,30,100,500$. At fixed active speed (across each row), increasing $\tilde{D}_{r}$ has a nuanced impact on the distribution of non-affinity. At high enough active speed, e.g., $\Lambda=6$, extremely large persistence ($\tilde{D}_{r}=0.01$) keeps the distribution unimodal, whereas producing a tail towards higher $\chi$ over an intermediate range of persistence. While persistence helps to produce bimodality, too large a persistence can produce jamming, suppressing the large $\chi$ fluctuations. Beyond this regime, progressive decrease in persistence suppresses the large-$\chi$ tail, and the distributions take unimodal form at high $\tilde{D}_{r}$.
Conversely, at fixed $\tilde{D}_{r}$ (down each column), increasing $\Lambda$ shifts probability weight toward larger $\chi$ and broadens $P(\chi)$. For low rotational diffusivity, $\tilde{D}_{r}<0.25$, this trend culminates in a clear bimodality at large $\Lambda$, reflecting the coexistence of a dominant low-$\chi$ population with a secondary high-$\chi$ contribution arising from rare, large non-affine rearrangements. Near $\tilde{D}_{r}=0.25$, the distributions show a strong enhancement of typical $\chi$, marked by a rightward shift and broadening of the main peak, signaling a crossover to the unimodal, tail-suppressed regime observed at larger $\tilde{D}_{r}$.
Appendix E Obtaining shear modulus
@@
Figure 14: (a) Shear stress, $\sigma_{xy}^{tot}$, vs. shear strain, $\mathcal{E}_{xy}=\gamma$, at $\Lambda=7$ for 24 ensembles without significant non-linearity or hysteresis. Different colors represent individual ensembles, with the black line showing the linear fit, where the slope gives the shear modulus, $\langle{\cal G}\rangle=105.7\pm 0.679$ in units of $\epsilon/\sigma^{2}$. (b) Linear response trajectories and (c) non-linear response trajectories exhibiting significant hysteresis and nonlinearity. Non-linear trajectories are excluded from shear modulus calculation. Raw data are shown as semi-transparent circular markers, with opaque square markers representing averages over 100 consecutive values of $\gamma$ and $\sigma^{\mathrm{tot}}_{xy}$. Markers are color-coded by shear direction: forward shear $0\rightarrow\gamma_{m}$ in red, backward shear $\gamma_{m}\rightarrow-\gamma_{m}$ in green, and $\gamma_{m}\rightarrow 0$ in blue. (d) Fraction of linear ensembles, $n_{\mathrm{lin}}/n_{\mathrm{tot}}$, as a function of activity $\Lambda$.
@@
${\cal E}_{xy}\Rightarrow 0\;\rightarrow\;+{\cal E}_{xy}^{\max}\;\rightarrow\;-{\cal E}_{xy}^{\max}\;\rightarrow\;0,$
at a strain rate $\dot{{\cal E}}=10^{-2}D_{r}$ and ${\cal E}_{xy}^{\max}=0.035$. The shear modulus is a linear response coefficient; thus the chosen ${\cal E}_{xy}^{\max}$ is to ensure the system response stays within the linear regime for accurate shear modulus extraction.
We perform the calculations at $\tilde{D}_{t}=1.0$ and $\tilde{D}_{r}=3.0$. For $\Lambda\leq 6$, all four initial configurations exhibit linear stress-strain behavior. For $\Lambda>6$, non-linear deviations appear. We generate $n_{\mathrm{tot}}\approx 100$ initial configurations, discarding those with nonlinearity for shear modulus analysis, which becomes more common at higher $\Lambda$. The shear modulus is then computed by averaging the linear response data from the remaining $n_{\mathrm{lin}}$ hysteresis-free ensembles, as shown in Fig. 14(a). Examples of linear and non-linear ensembles are shown in Figs. 14(b) and (c).